OpenStudy (anonymous):

Solve the equation: 7x^2 = -91

OpenStudy (anonymous):

7x^2 = -91 x^2 = -91/7 x^2 = -13 x^2 = +/- root(-13)

OpenStudy (anonymous):

\[ \left\{\left\{x\to -i \sqrt{13}\right\},\left\{x\to i \sqrt{13}\right\}\right\} \]

OpenStudy (anonymous):

what's the i before the 13?

OpenStudy (anonymous):

\[ \sqrt{-1} = i \] Read more here: http://en.wikipedia.org/wiki/Complex_number